Integration of sin^6x/cos^8x dx 
Please solve this question fast 

Grade:12

1 Answers

Aditya Gupta
5 years ago
this is very ez to do.
sin^6x/cos^8x= (sin^6x/cos^6x)*sec^2x= tan^6x*sec^2x
now substitute tanx=y or dy= sec^2x dx
so that ∫sin^6x/cos^8x dx = ∫tan^6x*sec^2x dx= ∫y^6 dy= y^7/7+C
tan^7x/7 + C